OpenStudy (anonymous):

Does anybody know how to solve rational equations with polynomial denominators?

OpenStudy (anonymous):

if so comment and i will post the question ^^

OpenStudy (yuki):

I do

OpenStudy (yuki):

I helped someone with that so I will guid you there

OpenStudy (anonymous):

Alrighty thx!! :) the problem is: x = 1 +1 2x+4 x+2

OpenStudy (anonymous):

i dont know how to do the lines underneath ^^

OpenStudy (yuki):

it's in clark's problem the idea is there, so if you want to know the steps they are all there

OpenStudy (anonymous):

Alrighty!!! can u link me maybe?

OpenStudy (yuki):

actually, yours is an equation. not an inequality, so nvm

OpenStudy (yuki):

uh... I just noticed this but what did I do ? lol there are so many postings... creepy

OpenStudy (anonymous):

Wow!! that looks like exactely what i need!!! :D thanks a bunch!!! :) Totally giving u medals lol :)

OpenStudy (yuki):

wait up 1<3, yours is different and much easier so stay here and listen

OpenStudy (anonymous):

alright im listening!! :)

OpenStudy (yuki):

\[{x \over 2x +4 } = {1 \over x+2 } +1 \]

OpenStudy (yuki):

right ?

OpenStudy (anonymous):

You can also watch Sal on khanacademy: http://www.khanacademy.org/v/solving-rational-equations-1?p=Algebra%20I%20Worked%20Examples

OpenStudy (anonymous):

Alright u guys have helped me massively!! :) Now i gtg and do more work :P might not be on later depends on if i need help or not or if i have time ^_^ lol

OpenStudy (yuki):

the first thing you want to do here is to find out the common denominator for all terms you have,

OpenStudy (anonymous):

gl!

OpenStudy (anonymous):

Thanks :)

OpenStudy (yuki):

I'll leave the steps so that you can look at it later ok ?

OpenStudy (anonymous):

Ahww thank you that is extremely helpful of u =3

OpenStudy (anonymous):

cuz i k i will definitely look at the steps ^_^

OpenStudy (yuki):

the LCD is 2x+4, so if you multiply that on both sides you will get \[x = 2 + 2x+4\]

OpenStudy (yuki):

if you solve that x = -6

OpenStudy (yuki):

the only thing you have to be careful is to make sure that when you start the problem, you identify the "extraneous solutions' basically the x's that makes any denominator in our eqn =0 since x = -2 is such number, it will never be the answer

OpenStudy (yuki):

since we did not deduce that x = -2, we just ignore it, but if x =-2 was what we deduced, then there is no solution. okay, hope you do well.

OpenStudy (anonymous):

Thanks a BUNCH!!!! ur steps totally helped me :)